OpenStudy (anonymous):

A cheetah is running at a speed of 65mph while a gazelle is running at 47mph. If the gazelle is 1mi ahead of the cheetah, how far must the cheetah run to catch the gazelle? Please include an expression.

OpenStudy (anonymous):

How do you take a screen shot help please

OpenStudy (perl):

lets use algebra let c(t), g(t) be distances of cheetah c(t) = 65*t g(t) = 47*t + 1

OpenStudy (anonymous):

I aggree

OpenStudy (perl):

then solve c(t) = g(t)

OpenStudy (perl):

65t = 47t + 1 18t = 1 t = 1/18

OpenStudy (anonymous):

Yay! you saved mah life, thanks :)

OpenStudy (perl):



OpenStudy (perl):